11. The common stock of Auto Deliveries sells for $25 a share. The stock is

expected to pay $1.10 per share next year when the annual dividend is distributed. The

firm has established a pattern of increasing its dividends by 3 percent annually and

expects to continue doing so. What is the dividend yield? What is the market rate of

return on this stock?

Dividend yield =4.4%

Market rate of return = 7.4%

Please Show All Work
Answer #2

Market rate of return on this stock = D1 / P0 + g

Market rate of return on this stock = $1.10 / $25 + 0.03

Market rate of return on this stock = 7.4%
Answer #1

Dividend Yield = Expected Dividend/Current Price of Share

Dividend Yield = $1.10/$25 = 4.40%

Based on constant growth dividend discount model,

Div Po=- Po Price of Stock Divi = Estimated Dividends for Next Period r = Required Rate of Return g Grouth Rate

1.10 r-0,03 25-

r - 0.03 = 0.044

Rate of Return, r = 7.4%
Answer #3

Given :


P0 = 25 ($)

D1 = 1.10 ($)

g = 3 % = 0.030



Now,


Dividend yield = D1/P0 

= 1.10/25 

= 0.044 

= 4.40% (ANSWER).


Market rate of return, r 

= Dividend yield + g

= 0.044 + 0.030

= 0.074

= 7.4 % (ANSWER)



answered by: Tulsiram Garg